# Index growth not imputable to topology

### v1

@@ 1 Introduction
The analysis of Hsiang’s [Hsiang1983] minimal hypersurfaces in $\mathbb{S}^{4}$ is the main object of Section 3. These hypersurfaces can be parametrized by the number of intersections with the “Clifford football” in $\mathbb{S}^{4}$ (as defined in Section 3). If we let $H_{m}$ denote the hypersurface with $m$ tori of intersection, then one can prove that the index grows at least linearly with $m$, with explicit estimates (see Theorem 3.11). In particular, in the case when the integer $m$ is odd, our result refers to the sequence of minimal hyperspheres produced by Hsiang to disporove Chern’s spherical Bernstein conjecture.
@@ Acknowledgments.
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 947923). The research of M. S. was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2044 – 390685587, Mathematics Münster: Dynamics–Geometry–Structure, and the Collaborative Research Centre CRC 1442, Geometry: Deformations and Rigidity.
@@ 2 Index estimates via partitioning methods for general group actions
Now let $G$ be a compact Lie group acting on $M$. We assume that the action $G\times M\ni(g,p)\mapsto g.p\in M$ is smooth and via isometries of $M$. We further assume that $G$ preserves (setwise) each of the sets $\Omega$, $\partial_{\mathrm{D}}\Omega$, $\partial_{\mathrm{N}}\Omega$, and $\partial_{\mathrm{R}}\Omega$. Additionally we let let $\sigma\colon G\to\operatorname{U}(1)$ be a Lie group homomorphism from $G$ to the unitary group on $\mathbb{C}$.
@@
Suppose (for this example) that $M$ is a two-sided embedded hypersurface with a global unit normal $\nu$ in some ambient Riemannian manifold $N$, suppose that $G$ is a subgroup of the isometry group of $N$ each of whose elements preserves $M$ as a set, and define $\sigma=\operatorname{sgn}_{\nu}\colon G\to\operatorname{O}(1)<\operatorname{U}(1)$ by
@@
Let $M$ be a smooth compact connected Riemannian manifold with (possibly empty boundary). Let $G$ be a compact Lie group acting smoothly on $M$ by isometries, and let $\sigma\colon G\to\operatorname{U}(1)$ be a homomorphism of Lie groups. Let $q,r\colon M\to\mathbb{R}$ be smooth functions invariant under $G$. Let $\Omega\subset M$ be a connected open subset of $M$, and let $\Omega_{1},\ldots,\Omega_{n}\subset\Omega$ be pairwise disjoint open subsets of $M$ whose closures cover $\Omega$. Assume further that each of the preceding $n+1$ open sets has Lipschitz boundary (possibly with $\partial\Omega=\emptyset$) and is preserved as a set by every element of $G$. Let $\partial_{\mathrm{D}}\Omega,\partial_{\mathrm{N}}\Omega,\partial_{\mathrm{R}}\Omega\subseteq\partial\Omega$ be pairwise disjoint open subsets of $\partial\Omega$ whose closures cover $\partial\Omega$ and each of which is preserved as a set by every element of $G$.
@@
Let $M$ be a smooth compact connected Riemannian manifold with boundary $\partial M$ and with interior $\mathring{M}$. Let $G$ be a compact Lie group acting smoothly on $M$ by isometries, and let $\sigma\colon G\to\operatorname{U}(1)$ be a homomorphism of Lie groups. Let $q,r\colon M\to\mathbb{R}$ be smooth functions invariant under $G$. Define
@@
as in (1), and let $\lambda\in\mathbb{R}$ be an eigenvalue of $T_{\mathrm{D}}$ having a $(G,\sigma)$-invariant eigenfunction. Then, recalling (2) for the definition of the following spaces of functions, we have
@@
By the commutativity of $\pi_{T_{\mathrm{R}}}^{\leq\lambda}$ and $\pi_{G,\sigma}$ we have
@@
It suffices therefore to check that the restriction of $\pi_{T_{\mathrm{R}}}^{\leq\lambda}$ to $E_{G,\sigma}^{<\lambda}(T_{\mathrm{D}})$ is injective. Suppose then that $u\in E_{G,\sigma}^{\leq\lambda}(T_{\mathrm{D}})$ is orthogonal in $L^{2}(\mathring{M})$ to $E^{<\lambda}_{G,\sigma}(T_{\mathrm{R}})$, and so in fact orthogonal to all of $E^{<\lambda}(T_{\mathrm{R}})$ (given that $u$ itself is $(G,\sigma)$-invariant, so orthogonal to the orthogonal complement of $\pi_{G,\sigma}(L^{2}(\mathring{M}))$). Then $u\in E^{\geq\lambda}(T_{\mathrm{R}})$, so
@@ 3 Applications to closed minimal hypersurfaces in round spheres
To put things in context, let us start with some recollections about what is known on closed (embedded) minimal hypersurfaces in the round four-dimensional sphere $\mathbb{S}^{4}$. Zhou [Zhou20] proved in 2020 the so-called multiplicity one conjecture (crucially relying on the earlier work of many, and primarily on the rich theory developed by Marques and Neves), which implies that $\mathbb{S}^{4}$ contains a sequence $\{\Sigma_{i}\}_{i\in\mathbb{N}}$ of closed minimal hypersurfaces, having area (volume) diverging as $i\to\infty$ precisely as prescribed by the Weyl law of Liokumovich–Marques–Neves [LioMarNev18], but with only limited topological control (if any) and no effective lower bounds on the Morse index of such hypersurfaces. That said, to the best of our knowledge, until about 1990 only two explicit examples were known (namely: the equatorial spheres $\mathbb{S}^{3}$ that have Morse index equal to 1, and the Clifford-type products $\mathbb{S}^{2}(\sqrt{2/3})\times\mathbb{S}^{1}(\sqrt{1/3})$ that have Morse index equal to 6) plus the two associated equivariant families constructed by Hsiang in [Hsiang1983], to be described below. Then, Solomon [Sol90] constructed an isomparametric minimal hypersurface (topologically a quotient of $\operatorname{SO}(3)$ by a suitable action of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$) and proved its Morse index to equal 20; this seems to be the only case, besides equatorial hyperspheres and Clifford-type hypersurfaces, where this invariant has been computed explicitly. More recently, the first and second author of the present paper proved in [CarSch21] the existence of one (conjecturally unique) minimal hypertorus in $\mathbb{S}^{4}$ (together with infinitely many immersed ones); the problem of computing its Morse index is open, but we shall note here that – as a corollary of [PerdomoLowIndex, Theorem 4.3] – such a hypertorus has Morse index at least $8$; we will get back to this matter later in this section, see in particular Remark 3.13 below.
@@
It is well-known (see for example Corollary 2.2 in [Sav10]) that for any two-sided minimal hypersurface $\Sigma\subset\mathbb{S}^{n+1}\subset\mathbb{R}^{n+2}$, the normal component (to $\Sigma$ in $\mathbb{S}^{4}$) of any translational Killing field on $\mathbb{R}^{n+2}$ is always (when nontrivial) an eigenfunction of eigenvalue $-n$ for the Jacobi operator $J$ of $\Sigma$, and the component functions of the unit normal are linearly independent whenever $\Sigma$ is not totally geodesic. However, to be self-contained we note how both claims can be easily seen by considering the cone $\widehat{\Sigma}$, centered at the origin, over $\Sigma$ in $\mathbb{R}^{n+2}$, as follows.
@@
taking values in $[-\frac{\pi}{4},\frac{\pi}{4}]$; thus $\vartheta$ is a normalized signed distance within the sphere of constant $x^{5}$ to $F$. Then $\vartheta$ is smooth with nowhere vanishing gradient away from $\{x^{3}=x^{4}=0\}\cup\{x^{1}=x^{2}=0\}$; note that the gradient of (the extension of) the function $\vartheta=\vartheta(x^{1},x^{2},x^{3},x^{4},x^{5})$ on $\mathbb{R}^{5}\setminus\left\{x^{1}=x^{2}=x^{3}=x^{4}=0\right\}$ has unit length and vanishing fifth component wherever it is defined. Furthermore, $\vartheta$ is constant (or undefined) on each orbit of $G$, as well as on each orbit of the projection $\partial_{5}^{\top}$ of the Killing field $\partial_{5}$ onto $\mathbb{S}^{4}$.
@@
The $G$-invariance of $H_{m}$ forces $\nu_{m}$ to be proportional to $\partial_{5}^{\top}$ on the two circular orbits of $G$ in $H_{m}$. Thus these circles are disjoint from the nodal set of $\nu^{5}$. Elsewhere $\vartheta|_{H_{m}}$ is smooth, and a point $p\in H_{m}$ is a critical point of $\vartheta|_{H_{m}}$ if and only if $\nu_{m}|_{p}$ is equal (possibly modulo sign) to $\nabla\vartheta|_{p}$, but $\nabla\vartheta|_{p}$ is orthogonal (along the toroidal $G$-orbit through $p$) to $\partial_{5}^{\top}|_{p}$, establishing the claim. ∎
@@
By Lemma 3.9 we have a partition of $H_{m}$ into $m-1$ domains (all of whose boundary components are smooth tori), for each of which $-3$ is a second Neumann eigenvalue, with multiplicity one. Thus, exploiting item (ii) of Proposition 2.2, we have one domain accounting for a $+2$ contribution, and exactly $m-2$ domains accounting for a $+1$ contribution, which gives the claim. ∎
@@
Firstly, it follows from Proposition 2.2 and Lemma 3.8 that the Jacobi operator $J_{m}$ acting on the space of $G$-invariant functions on $H_{m}$ has at least $m-1$ eingenvalues strictly less than $-3$, plus the eigenvalue $-3$ with multiplicity at least one (the corresponding eingenfunction being $\nu_{m}^{5}$); this fact together with the upper bound given by Proposition 3.10 proves (i). In particular, the $G$-equivariant Morse index of $H_{m}$ is at least $m$. If we lift the equivariance constraint, we still have a subspace of dimension at least $m-1$ corresponding to the eigenvalues strictly less than $-3$ plus, in direct sum, a five-dimensional subspace spanned by eigenfunctions with eigenvalue exactly equal to $-3$ (because of Lemma 3.1). The conclusion of (ii) follows at once. ∎
@@ 4 Applications to free boundary minimal hypersurfaces in Euclidean balls
Now, consider instead a critical point $y$ on the $\left\{\rho=0\right\}$ orbit. Here $A$ has vector-valued second fundamental form $\operatorname{II}_{y}(v,v)$ directed radially inward for $v$ tangent to the orbit (as for the toroidal orbits considered above). The two-dimensional subspace of $\operatorname{T}_{y}A$ orthogonal to $v$ is invariant under the subgroup of $G$ that fixes $y$ (a copy of $\operatorname{O}(2)$). By the symmetry we have $\operatorname{II}_{y}(u,u)=\operatorname{II}_{y}(w,w)$ for any unit vectors $u$ and $w$ in this subspace. By the minimality of $A$ we must then have $\operatorname{II}_{y}(u,u)$ directed radially outward for every such $u$, and so also in this case $\lvert x\rvert|_{A}$ is strictly increasing as one moves away from the point $y$ in directions orthogonal to the $G$-orbit.
@@
$\{x\in A\,:\ x^{k}\nu_{k}(x)=0\}=\bigcup_{m=1}^{\infty}\{x\in A\,:\ \lvert x\rvert=r_{m}\},$
@@
The $G$-invariant Jacobi field $x^{k}\nu_{k}$ has $\ell$ nodal domains and satisfies the homogeneous Dirichlet condition on $\partial A_{\ell}$. It then follows (from item (i) of Proposition 2.2, applied for $q$ equal to the squared length of the second fundamental form of $A_{\ell}$) that the Jacobi operator for $A_{\ell}$ acting on $G$-invariant functions (equivalently: $G$-equivariant functions, cf. Example 2.1) subject to the homogeneous Dirichlet condition has at least $\ell$ nonpositive eigenvalues (counted with multiplicity). At that stage the conclusion comes by simply appealing to Proposition 2.3, which allows to compare the preceding index bound with the geometrically relevant one where oblique boundary conditions are imposed (for $r\equiv 1$). ∎

### v2

@@ 1 Introduction
The analysis of Hsiang’s [Hsiang1983] minimal hypersurfaces in $\mathbb{S}^{4}$ is the main object of Section 3. These hypersurfaces can be parametrized by the number of connected components in the intersection with the “Clifford football” in $\mathbb{S}^{4}$ (as defined in Section 3). If we let $H_{m}$ denote the hypersurface with $m$ tori of intersection, then one can prove that the index grows at least linearly with $m$, with explicit estimates (see Theorem 3.11). In particular, in the case when the integer $m$ is odd, our result refers to the sequence of minimal hyperspheres produced by Hsiang to disporove Chern’s spherical Bernstein conjecture.
@@ Acknowledgments.
The authors thank the anonymous referee for their valuable suggestions. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 947923). The research of M. S. was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2044 – 390685587, Mathematics Münster: Dynamics–Geometry–Structure, and the Collaborative Research Centre CRC 1442, Geometry: Deformations and Rigidity.
@@ 2 Index estimates via partitioning methods for general group actions
Now let $G$ be a compact Lie group acting on $M$. We assume that the action $G\times M\ni(g,p)\mapsto g.p\in M$ is smooth and via isometries of $M$ and that $q(g.p)=q(p)$ and $r(g.p)=r(p)$ for all $g\in G$ and $p\in M$. We further assume that $G$ preserves (setwise) each of the sets $\Omega$, $\partial_{\mathrm{D}}\Omega$, $\partial_{\mathrm{N}}\Omega$, and $\partial_{\mathrm{R}}\Omega$. Additionally we let $\sigma\colon G\to\operatorname{U}(1)$ be a Lie group homomorphism from $G$ to the unitary group on $\mathbb{C}$.
@@
Suppose (for this example) that $M$ is a two-sided, connected embedded hypersurface with a global unit normal $\nu$ in some ambient Riemannian manifold $N$, suppose that $G$ is a subgroup of the isometry group of $N$ each of whose elements preserves $M$ as a set, and define $\sigma=\operatorname{sgn}_{\nu}\colon G\to\operatorname{O}(1)<\operatorname{U}(1)$ by
@@
Let $M$ be a smooth compact connected Riemannian manifold with (possibly empty) boundary. Let $G$ be a compact Lie group acting smoothly on $M$ by isometries, and let $\sigma\colon G\to\operatorname{U}(1)$ be a homomorphism of Lie groups. Let $q,r\colon M\to\mathbb{R}$ be smooth functions invariant under $G$. Let $\Omega\subset M$ be a connected open subset of $M$, and let $\Omega_{1},\ldots,\Omega_{n}\subset\Omega$ be pairwise disjoint open subsets of $M$ whose closures cover $\Omega$. Assume further that each of the preceding $n+1$ open sets has Lipschitz boundary (possibly with $\partial\Omega=\emptyset$) and is preserved as a set by every element of $G$. Let $\partial_{\mathrm{D}}\Omega,\partial_{\mathrm{N}}\Omega,\partial_{\mathrm{R}}\Omega\subseteq\partial\Omega$ be pairwise disjoint open subsets of $\partial\Omega$ whose closures cover $\partial\Omega$ and each of which is preserved as a set by every element of $G$.
@@
Let $M$ be a smooth compact connected Riemannian manifold with boundary $\partial M\neq\emptyset$ and with interior $\mathring{M}$. Let $G$ be a compact Lie group acting smoothly on $M$ by isometries, and let $\sigma\colon G\to\operatorname{U}(1)$ be a homomorphism of Lie groups. Let $q,r\colon M\to\mathbb{R}$ be smooth functions invariant under $G$. Define
@@
as in (1), and let $\lambda\in\mathbb{R}$. Then, recalling (2) for the definition of the following spaces of functions, we have
@@
By the commutativity of $\pi_{T_{\mathrm{R}}}^{<\lambda}$ and $\pi_{G,\sigma}$ we have
@@
It suffices therefore to check that the restriction of $\pi_{T_{\mathrm{R}}}^{<\lambda}$ to $E_{G,\sigma}^{\leq\lambda}(T_{\mathrm{D}})$ is injective. Suppose then that $u\in E_{G,\sigma}^{\leq\lambda}(T_{\mathrm{D}})$ is orthogonal in $L^{2}(\mathring{M})$ to $E^{<\lambda}_{G,\sigma}(T_{\mathrm{R}})$, and so in fact orthogonal to all of $E^{<\lambda}(T_{\mathrm{R}})$ (given that $u$ itself is $(G,\sigma)$-invariant, so orthogonal to the orthogonal complement of $\pi_{G,\sigma}(L^{2}(\mathring{M}))$). Then $u\in E^{\geq\lambda}(T_{\mathrm{R}})$, so
@@ 3 Applications to closed minimal hypersurfaces in round spheres
To put things in context, let us start with some recollections about what is known on closed (embedded) minimal hypersurfaces in the round four-dimensional sphere $\mathbb{S}^{4}$. Zhou [Zhou20] proved in 2020 the so-called multiplicity one conjecture (crucially relying on the earlier work of many, and primarily on the rich theory developed by Marques and Neves), which implies that $\mathbb{S}^{4}$ contains a sequence $\{\Sigma_{i}\}_{i\in\mathbb{N}}$ of closed minimal hypersurfaces, having area (volume) diverging as $i\to\infty$ precisely as prescribed by the Weyl law of Liokumovich–Marques–Neves [LioMarNev18], but with only limited topological control (if any) and no effective lower bounds on the Morse index of such hypersurfaces. That said, to the best of our knowledge, only very few “explicit” examples are known. Among these, we first mention the equatorial spheres $\mathbb{S}^{3}$ that have Morse index equal to 1, and the Clifford-type products $\mathbb{S}^{2}(\sqrt{2/3})\times\mathbb{S}^{1}(\sqrt{1/3})$ (that have Morse index equal to 6) plus the two associated equivariant families constructed by Hsiang in [Hsiang1983], to be described below. To move beyond such instances, we shall recall Cartan’s work [Cartan1939] on the construction and classification of cubic isoparametric minimal hypersurfaces in spheres; it turns out that there is precisely one in $\mathbb{S}^{4}$, and that is topologically a quotient of $\operatorname{SO}(3)$ by a suitable action of $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. Solomon then proved in [Sol90] that its Morse index is equal to $20$: this seems to be the only case, besides equatorial hyperspheres and Clifford-type hypersurfaces, where this invariant has been computed explicitly. More recently, the first and second author of the present paper proved in [CarSch21] the existence of one (conjecturally unique) minimal hypertorus in $\mathbb{S}^{4}$ (together with infinitely many immersed ones); the problem of computing its Morse index is open, but we shall note here that – as a corollary of [PerdomoLowIndex, Theorem 4.3] – such a hypertorus has Morse index at least $8$; we will get back to this matter later in this section, see in particular Remark 3.13 below.
@@
It is well-known (see for example Corollary 2.2 in [Sav10]) that for any two-sided minimal hypersurface $\Sigma\subset\mathbb{S}^{n+1}\subset\mathbb{R}^{n+2}$, the normal component (to $\Sigma$ in $\mathbb{S}^{n+1}$) of any translational Killing field on $\mathbb{R}^{n+2}$ is always (when nontrivial) an eigenfunction of eigenvalue $-n$ for the Jacobi operator $J$ of $\Sigma$, and the component functions of the unit normal are linearly independent whenever $\Sigma$ is not totally geodesic. However, to be self-contained we note how both claims can be easily seen by considering the cone $\widehat{\Sigma}$, centered at the origin, over $\Sigma$ in $\mathbb{R}^{n+2}$, as follows.
@@
taking values in $[-\frac{\pi}{4},\frac{\pi}{4}]$; thus $\vartheta$ is a normalized signed distance within the sphere of constant $x^{5}$ to $F$. Then $\vartheta$ is smooth with nowhere vanishing gradient away from $\{x^{3}=x^{4}=0\}\cup\{x^{1}=x^{2}=0\}$. Furthermore, $\vartheta$ is constant (or undefined) on each orbit of $G$, as well as on each orbit of the projection $\partial_{5}^{\top}$ of the Killing field $\partial_{5}$ onto $\mathbb{S}^{4}$.
@@
The $G$-invariance of $H_{m}$ forces $\nu_{m}$ to be proportional to $\partial_{5}^{\top}$ on the two circular orbits of $G$ in $H_{m}$. Thus these circles are disjoint from the nodal set of $\nu^{5}$. Elsewhere $\vartheta|_{H_{m}}$ is smooth, and a point $p\in H_{m}$ is a critical point of $\vartheta|_{H_{m}}$ if and only if $\nu_{m}|_{p}$ is proportional to $\nabla\vartheta|_{p}$, but $\nabla\vartheta|_{p}$ is orthogonal to $\partial_{5}^{\top}|_{p}$ as well as to the toroidal $G$-orbit through $p$, establishing the claim. ∎
@@
By Lemma 3.9 we have a partition of $H_{m}$ into $m-1$ domains (all of whose boundary components are smooth tori), for each of which $-3$ is a second Neumann eigenvalue, with multiplicity one. (Note that when $m=2$ there is no actual partition, i. e. we are just considering $H_{2}$ itself and thus no boundary conditions come into play.) Thus, exploiting item (ii) of Proposition 2.2, we have one domain accounting for a $+2$ contribution, and exactly $m-2$ domains accounting for a $+1$ contribution, which gives the claim. ∎
@@
According to Proposition 3.4, $J_{m}$ acting on the space of $G$-invariant functions on $H_{m}$ has at least $m-1$ eingenvalues strictly less than $-3$, plus the eigenvalue $-3$ with multiplicity at least one (the corresponding eingenfunction being $\nu_{m}^{5}$); this fact together with the upper bound given by Proposition 3.10 proves (i). In particular, the $G$-equivariant Morse index of $H_{m}$ is at least $m$. If we lift the equivariance constraint, we still have a subspace of dimension at least $m-1$ corresponding to the eigenvalues strictly less than $-3$ plus, in direct sum, a five-dimensional subspace spanned by eigenfunctions with eigenvalue exactly equal to $-3$ (because of Lemma 3.1). The conclusion of (ii) follows at once. ∎
@@ 4 Applications to free boundary minimal hypersurfaces in Euclidean balls
Now, consider instead a critical point $y$ on the $\left\{\rho=0\right\}$ orbit. Here $A$ has vector-valued second fundamental form $\operatorname{II}_{y}(v,v)$ directed radially inward for $v$ tangent to the orbit (as for the toroidal orbits considered above). The two-dimensional subspace of $T_{y}A$ orthogonal to $v$ is invariant under the subgroup of $G$ that fixes $y$ (a copy of $\operatorname{O}(2)$). By the symmetry we have $\operatorname{II}_{y}(u,u)=\operatorname{II}_{y}(w,w)$ for any unit vectors $u$ and $w$ in this subspace. By the minimality of $A$ we must then have $\operatorname{II}_{y}(u,u)$ directed radially outward for every such $u$, and so also in this case $\lvert x\rvert|_{A}$ is strictly increasing as one moves away from the point $y$ in directions orthogonal to the $G$-orbit.
@@
$\bigl{\{}x\in A\,:\ x^{k}\nu_{k}(x)=0\bigr{\}}=\bigcup_{m=1}^{\infty}\{x\in A\,:\ \lvert x\rvert=r_{m}\},$
@@
The $G$-invariant Jacobi field $x^{k}\nu_{k}$ has $\ell$ nodal domains and satisfies the homogeneous Dirichlet condition on $\partial A_{\ell}$. It then follows (from item (i) of Proposition 2.2, applied for $q$ equal to the squared length of the second fundamental form of $A_{\ell}$) that the Jacobi operator for $A_{\ell}$ acting on $G$-invariant functions (equivalently: $G$-equivariant functions, cf. Example 2.1) subject to the homogeneous Dirichlet condition has at least $\ell-1$ strictly negative eigenvalues (counted with multiplicity), and so, including $x^{k}\nu_{k}$ itself, at least $\ell$ nonpositive eigenvalues. At this stage the conclusion comes by simply appealing to Proposition 2.3, which allows us to compare the preceding bound with the geometrically relevant one where oblique boundary conditions are imposed (for $r\equiv 1$). ∎